15-th Conference of the International Linear Algebra Society

April 25, 2008

Professor Kenzi Satô

Tamagawa University Tokyo Japan

Dear Professor Satô:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “The algebraic relations of curvatures of PL manifolds” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Zeyad Al Zhour

Zarqa Private University Zarqa Jordan

Dear Professor Al Zhour:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Matrix Results on Weighted Drazin Inverse and Some Applications (with Adem Kilicman)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor María Isabel Bueno Cachadina

The University of California at Santa Barbara Santa Barbara USA

Dear Professor Bueno Cachadina:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Algorithms for computing the Geronimus Transformation” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Nieves Castro-González

Universidad Politécnica de Madrid Madrid Spain

Dear Professor Castro-González:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Representations for the generalized Drazin inverse of additive perturbations (with M.F. Martínez-Serrano)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Vanesa Cortés

Universidad de Zaragoza Zaragoza Spain

Dear Professor Cortés:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Some properties of the class sign regular matrices and its subclasses (with J.M. Peña)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Henrique F. da Cruz

U.B.I Covilhã Portugal

Dear Professor da Cruz:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On the matrices that preserve the value of the immanant of the upper triangular matrices (with Rosário Fernandes)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Fernando De Terán

Universidad Carlos III de Madrid Leganés Spain

Dear Professor De Terán:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Linearizations of Singular Matrix Polynomials and the Recovery of Minimal Indices (with Froilán M. Dopico and D. Steven Mackey)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor María Elena Domínguez

Universidad Politécnica de Madrid Madrid Spain

Dear Professor Domínguez:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “General solution of certain matrix equations arising in filter design applications” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Ajda Fošner

Institute of Mathematics, Physics and Mechanics Ljubljana Slovenia

Dear Professor Fošner:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Commutativity preserving maps on real matrices” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Maria T. Gassó

Inst.Mat.Mult. Universidad Politécnica Valencia Valencia Spain

Dear Professor Gassó:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “The class of Inverse-Positive matrix with chekecboard pattern (with Manuel F. Abad, and Juan R. Torregrosa)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor José Ramón Guzmán

Instituto de Investigaciones Económicas. UNAM. México México

Dear Professor Guzmán:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Reduction of an Itos diffusion input output model for the determination of square mean stability” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Hosoo Lee

Kyungpook National University Daegu Korea

Dear Professor Lee:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Contractions and nonlinear matrix equations on positive definite cones (with Yongdo Lim)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor José-Javier Martínez

Universidad de Alcalá Alcalá de Henares Spain

Dear Professor Martínez:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Polynomial regression in the Bernstein basis (with Ana Marco)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Marta Peña

Universitat Politecnica de Catalunya Barcelona Spain

Dear Professor Peña:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Perturbations preserving conditioned invariant subspaces (with A. Compta and J. Ferrer)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Cecília Perdigão

Faculdade de Cincias e Tecnologia-UNL Lisboa Portugal

Dear Professor Perdigão:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “On the equivalence class graph (with Rosário Fernandes)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Isabel Teixeira Matos

Centro de Estruturas Lineares e Combinatórias (CELC) Lisboa Portugal

Dear Professor Teixeira Matos:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “A Completion Problem over the Field of Real Numbers” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Xaab Nop Vargas Vásquez

ICYTDF México México

Dear Professor Vargas Vásquez:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “Students difficulties with concept of vector space from point of view of APOS Theory” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx

April 25, 2008

Professor Juan Alberto Rodríguez Velázquez

Universitat Rovira i Virgili Tarragona España

Dear Professor Rodríguez Velázquez:

The Organizing Committee of the 15-th Conference of the International Linear Algebra Society has accepted your paper “The Laplacian Spectrum of Hypergraphs (with Aida Kamisalic)” to be presented in the Contributed Papers Session.

You will have twenty minutes for your presentation and five minutes for questions and comments. We will inform you about the date and time of your talk around May 5.

We remind you that the 15-th Conference will be held in Cancún, México, from June 16 to June 20, 2008.

We are certain that your lecture will contribute greatly towards the success of our conference and will be highly appreciated by all the participants.

Cordially Dr. Luis Verde-Star

Chair of the Organizing Committee, ILAS Cancún 2008

Departamento de Matemáticas, UAM Iztapalapa

Apartado 55-534, Mexico DF 09340, Mexico

Tel. +52 5558044961, Fax +52 5558044660, verde@star.izt.uam.mx